OpenStudy (anonymous):

Y varies indirectly as twice x. When x = 1, y = 3. Find y when x = 4.

OpenStudy (jamesj):

'indirectly' is not a word working mathematicians use, but I am pretty sure this means that y = mx + c for some constants m and c. Now saying x "varies ... as twice x" I take to mean m = 2. That being the case, if x = 1 implies y = 3. Then y = mx + c 3 = 2 times 1 + c 3 = 2 + c c = 1 Hence y = 2x + 1 Thus, when x = 4, y = ....

OpenStudy (jamesj):

( @agd: you wrote down an equation for "y varies INVERSELY with x" )

OpenStudy (anonymous):

I don't know what indirectly means :-(

OpenStudy (anonymous):

thats alright....neither do i :(

OpenStudy (anonymous):

xy = a constant; 1*3=3 so all xy products must equal 3. 4*x=3 so x=3/4